OpenStudy (anonymous):

Find the quotient. write in simplest form:

OpenStudy (anonymous):

\[\frac{ q }{ 7 } \div \frac{ q }{ 26 }\]

OpenStudy (anonymous):

argh the fraction isn't working

OpenStudy (anonymous):

q/7 div q/26

OpenStudy (anonymous):

awww sorry katt i don't know this:(( i will give u a medal tho bc you rock!!

OpenStudy (shamil98):

\[\frac{ q }{ 7 } \times \frac{ 26 }{ q }\]

OpenStudy (shamil98):

when you divide fractions you multiply by the reciprocal

OpenStudy (anonymous):

ohh okay thanks @shamil98 and its okay twinz! haha thanks!

OpenStudy (shamil98):

\[\frac{ \cancel q }{ 7 } \times \frac{ 26 }{ \cancel q } = \frac{ 26 }{ 7 }\]

OpenStudy (anonymous):

thanks!!